Solve for b2 A=1/2h(b1+b2)

Solve for b2 A=1/2h(b1+b2)
Answer #1
Solve for b2 A=1/2h(b1+b2) : A = {{{1/2}}}h(b1 + b2) : Multiply both side by 2, and you have: 2A = h(b1 + b2) : Divide both sides by h, resulting in: {{{(2A)/h}}} = b1 + b2 : Subtract b1 from both sides: {{{(2A)/h}}}- b1 = b2
answered by: kong